The following items are independent. Assume that the original transactions have been recorded correctly or as described. Assume a December 31 year-end unless otherwise noted.

  1. Prepaid insurance had a debit balance of $16,800 at the beginning of January. This represents the remaining 14 months in an insurance policy that was purchased in a prior year. On 1 April of the current year, a 30-month policy was bought for $46,600, which was debited to prepaid insurance. There were no other entries to the prepaid insurance account during the year. On 1 November, an 18-month policy was bought for $9,180. This policy was debited to insurance expense.
  2. Certain invoices had not been recorded at 31 December: a $6,900 bill for power, a $700 phone bill, and a repair bill for $1,250. All these bills are due in January.
  3. Salaries payable has a balance of $14,700, unadjusted from the last year-end. There are 14 employees. They were paid up to Friday, 27 December. There were two additional working days prior to the year-end. Three employees earn $700 per (five-day) week each, three earn $700 per week, and eight earn $700 per week.
  4. There are two notes payable outstanding. The first is an $1,000,000, 4% note, issued on 1 September; no interest has been paid to date. The second is a $786,900, 3.5% note issued on 1 April; six months’ interest was paid on 1 October. The notes payable themselves were properly recorded on issuance, but interest must be accrued to 31 December, the year-end.
  5. There is an unearned revenue account with a balance of $161,100. This includes a $11,100 security deposit from a tenant. This deposit will be returned when the tenant eventually vacates. The remaining $150,000 is six months’ rent received, on a lease beginning 1 December, on commercial property.
  6. Supplies inventory had a balance of $23,700 at the end of last year. During the year, $70,200 of supplies were purchased, and were debited directly to supplies expense. At year-end, an inventory count was conducted, and the balance was ascertained to be $24,800.
  7. Advertising expense of $95,300 includes a $6,800 payment for an advertisement that will run in January of next   1. Prepare adjusting journal entries to reflect the facts above.   2. Prepare reversing entries for adjusting journal entries that must be reversed.​​​​​​​
Answer #1

Adjusting Entries

Account Titles Debit Credit
Insurance Expense $          20,220 =16800/14*12+46600/30*9-9180/18*16
       Prepaid Insurance $          20,220
Utilities and Repair Expense $             8,850 =6900+700+1250
       Accounts Payable $            8,850
Salaries Payable $          10,780 =14700-(700/5*2*14)
       Salaries Expense $          10,780
Interest Expense $          20,219 =1000000*4%*4/12+786900*3.5%*3/12
       Interest Payable $          20,219
Unearned Revenue $          25,000 =150000/6
       Rent Revenue $          25,000
Supplies $             1,100 =70200-(23700+70200-24800)
       Supplies Expense $            1,100
Prepaid Advertising expense $             6,800
       Advertising Expense $            6,800

Reversing Entries

Account Titles Debit Credit
Accounts Payable $             8,850
        Utilities and Repair Expense $            8,850
Salaries Payable $             3,920
       Salaries Expense $            3,920
Interest Payable $          20,219
       Interest Expense $          20,219

If you have any query, kindly comment with your query and please mark thumbs up.